Solve systems of equation by graphing, then classify the system
9x-5y=-25
5y-9x=25

Answer by Alan3354(69443)

Solve systems of equation by graphing, then classify the system
9x-5y=-25
5y-9x=25
-----------
5y-9x=25 --> 9x-5y = -25
It's not a system, it's the same equation twice.
That makes it dependent.